iuytyuioiuytyuiop:

Given the function h(c)=8c2−41c, solve for h(c)=42. Give an exact answer; do not round. (Use a comma to separate multiple solutions.)

iuytyuioiuytyuiop:

@AZ

iuytyuioiuytyuiop:

az

iuytyuioiuytyuiop:

help? i need it

snowflake0531:

Can you use desmos? Or....?

iuytyuioiuytyuiop:

graphing calculator or scientific calculator

iuytyuioiuytyuiop:

or matrix calculator

snowflake0531:

I recommend desmos

iuytyuioiuytyuiop:

but desmos has different calculators

snowflake0531:

You can put in 8c^2 -41c and also y=41 The two points that get intersected are the points

iuytyuioiuytyuiop:

az left?!

snowflake0531:

am i not enough?

iuytyuioiuytyuiop:

your more!!!!!!!

snowflake0531:

*you're

iuytyuioiuytyuiop:

41 is the anser?

snowflake0531:

https://www.desmos.com/calculator/ytpfqxw2ys Look at the points of intersection

iuytyuioiuytyuiop:

i cnat see it

carmelle:

"cnat"

snowflake0531:

I suggest scrolling up in the graph

snowflake0531:

There are two points of intersection, one negative, one positive, look up

snowflake0531:

1 attachment
snowflake0531:

See those points of intersection? You can also click on them, and it'll tell you the exact coordinate

snowflake0531:

Or, we could also do this a different way, we can use the quadratic formula:))))) Do you want to do that

iuytyuioiuytyuiop:

-0.8, and 6

snowflake0531:

There's something after -0.8 It's -0.85

snowflake0531:

Do you want to redo this and do it using the quadratic formula? it's nice practice

iuytyuioiuytyuiop:

alright

snowflake0531:

Kay Do you know the quadratic formula?

iuytyuioiuytyuiop:

i only kno a small amount

snowflake0531:

1 attachment
snowflake0531:

We have 42 = 8c^2 -41c Because we want it all on one side, we do 8c^2 -41c -42

snowflake0531:

Yea?

iuytyuioiuytyuiop:

yep

snowflake0531:

Okay, so now we plug it in:)

snowflake0531:

Look at the screenshot, and remember the a, b, and c, are from ax^2 + bx+c, aka, 8c^2-41c-42

snowflake0531:

So, first part, -b What is b, what is -b?

iuytyuioiuytyuiop:

-0.85?

snowflake0531:

we have ax^2 + bx+ c 8c^2 -41c-42 a=8 b= -41 c= -42

iuytyuioiuytyuiop:

alright oh i see

snowflake0531:

So, plug it in, first step -b What is it

iuytyuioiuytyuiop:

-41

snowflake0531:

b=41 what is -b

iuytyuioiuytyuiop:

im not sure

snowflake0531:

What is -(-41)

iuytyuioiuytyuiop:

41

snowflake0531:

Okay next step sqrt(b^2 -4ac) What is b^2, remembering that b is -41

iuytyuioiuytyuiop:

-41^2

snowflake0531:

So what does that equal to

iuytyuioiuytyuiop:

-1681

snowflake0531:

._. Why would it still be negative

iuytyuioiuytyuiop:

1681

snowflake0531:

Okay, and what is 4ac 4 times a times c remembering that a is 8, c is -42

iuytyuioiuytyuiop:

4 * 8 * -42

snowflake0531:

So what does that equal to

iuytyuioiuytyuiop:

1344

snowflake0531:

Why did you get rid of the negative sign._.

iuytyuioiuytyuiop:

-1344

snowflake0531:

Kk, so then what is 1681-(-1344)

iuytyuioiuytyuiop:

3025

snowflake0531:

Okay, so now we got 41 +- sqrt3025 now we gotta do the denominator 2a a is 8 What is 2a

iuytyuioiuytyuiop:

16

snowflake0531:

Kay So it is (41 +- sqrt3025)/16 Can you tell me what sqrt 3025 is

iuytyuioiuytyuiop:

55

snowflake0531:

Okay, nowww we got to the part where we have to do two equations (41+55)/16 and 41-55)/16 Can you tell me what are the results of both

iuytyuioiuytyuiop:

96 and -14

snowflake0531:

They have to be divided by 16... What is 96/16 and -14/16

iuytyuioiuytyuiop:

6 and -0.875

snowflake0531:

There are your 2 answers! So, in the future, you can do it the graphing method or use the quadratic formula

iuytyuioiuytyuiop:

graphing method!!!!!!!!!!!!!

iuytyuioiuytyuiop:

ho do i graph this Given the function h(a)=9a2+46a, solve for h(a)=−5.

snowflake0531:

post a new question